OpenStudy (anonymous):

solve for n: squareroot(n+5)= n-1

OpenStudy (anonymous):

i got 4?

OpenStudy (anonymous):

n+5=n^2+1-2n =>n^2-3n-4 n=4,-1 but n=-1 is not possible...so n=4 only.

OpenStudy (anonymous):

squre both sides n+ 5) =(n -1)^2 n^2 - 2n + 1 -n - 5 = 0 n^2 -3n -4 = 0 x = 4 or -1

OpenStudy (anonymous):

why isn't -1 possible?

OpenStudy (anonymous):

squaring both sides u get n+5 = (n-1)^2 so n+5 = n^2 - 2n +1 0= n^2 - 2n -n +1 -5 0 = n^2 -3n -4 0 = n^2 -4n+1n-4 0=n(n-4)+1(n-4) 0=(n+1)(n-4) now this product can be zero if either (n+1) is zero or if (n-4) is zero if n+1 = 0 then n= -1 if n-4 =0 then n = 4 so u hv two value of n i.e -1 and 4

OpenStudy (anonymous):

yes - -1 is possible

OpenStudy (anonymous):

if we put n = -1 we get squareroot (-1+5) = -1-1 squareroot (4) = -2 this is a valid answer becoz (-2)^2 = 4 !!!!

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

so both n = 4 and n= -1 are valid @brackett what is your view on this ?????